The 1.4 m/s motion of a freely coasting tank car is momentarily stopped by a spring bumper comprised of two un-compressed springs arranged end-to-end as shown having spring constants kA = 6100 N/mm and kB = 19520 N/mm. The total compression at the instant the tank car stops is 0.38 m. Determine the mass of the tank car in kg.
Answer #1

net spring constant

1/k = 1/k1 + 1/k2

1/k = 1/19520 + 1 / 6100

k = 4647.619 N/m

using energy conservation

0.5 k x^2 = 0.5 m v^2

4647.619* 0.38^2 = m * 1.4^2

m = 342.406 kg

=====

Comment before rate in case any doubt, will reply for sure.. goodluck